# CRMAgent: A Multi-Agent LLM System for E-Commerce CRM Message Template Generation

### v1

@@ Abstract
CRMAgent: A Multi-Agent LLM System for E-Commerce
CRM Message Template Generation
Yinzhu Quan Georgia Institute of Technology Atlanta, GA 30332, USA yquan9@gatech.edu Xinrui Li ByteDance Inc. Seattle, WA 98004, USA ava.li@bytedance.com Ying Chen ByteDance Inc. San Jose, CA 95110, USA eric.ychen@bytedance.com
1 Introduction
Customer Relationship Management (CRM) Kumar and Reinartz (2018) messages serve as a primary lever for user retention and revenue generation on modern e-commerce platforms. Merchants routinely push promotional notifications that announce new products, highlight price drops, or deliver loyalty rewards to segmented audiences through app and messaging channels. Crafting effective templates remains a bottleneck because most merchants lack dedicated marketing expertise, work under tight time constraints, and therefore rely on generic system-generated copy. Many CRM messages consequently underperform, leaving substantial engagement and conversion potential untapped.
Recent work has shown the effectiveness of LLM agents for structured decision-making in operational domains such as logistics Ieva et al. (2025); Quan et al. (2025), supply chain management Li et al. (2023); Quan and Liu (2024b); Wang et al. (2025), and economic reasoning Quan and Liu (2024a); Guo and Yang (2024); Liu and Quan (2025a). To address this performance gap in CRM, we present CRMAgent (Figure 1), a multi-agent system built on large language models (LLMs) that automatically improves underperforming CRM message templates in e-commerce. Using historical engagement data, CRMAgent identifies low-performing messages, retrieves high-quality examples from similar contexts, and produces improved variants tailored to each merchant and audience segment group. The workflow is divided among four specialized agents: ContentAgent, RetrievalAgent, TemplateAgent, and EvaluateAgent, each responsible for a distinct subtask.
@@ 1 Introduction
2 Related Work
Recent multi-agent LLM systems adopt modular architectures to address structured generation tasks across diverse domains. For instance, Singh et al. (2024) design role-based agents for collaborative travel planning, while DocAgent Yang et al. (2025) structures code documentation into sequential reader–writer–verifier stages. MAxPrototyper Yuan et al. (2024) coordinates theme, text, and visual elements under a central controller for UI prototyping, and InvAgent Quan and Liu (2024b) demonstrates that agent specialization benefits inventory management in logistics. Similarly, CRMAgent applies a task-specialized multi-agent framework to CRM message rewriting, with agents dedicated to content analysis, retrieval, generation, and evaluation. Complementary to modularity, retrieval-augmented generation (RAG) improves LLM output by grounding in exemplar data or structured knowledge. Retrieve-and-Edit Hashimoto et al. (2018), PEARL Malik et al. (2024), and Chen et al. (2024) leverage retrieval to enhance preference modeling and factual accuracy. Recent work also explores temporal event modeling and retrieval Liu and Quan (2024, 2025b), combining LLMs with temporal point processes to unify event semantics and timing, while Liu et al. (2024) show that fine-tuning with retrieved exemplars improves personalization. Extending these directions, CRMAgent retrieves high-performing CRM templates based on audience, product, and voucher metadata, and generates tailored messages using marketing feedback.
3 Multi-Agent Collaboration for CRM Message Template Generation
@@ 3.1 ContentAgent
3.2 RetrievalAgent
RetrievalAgent retrieves high-quality CRM message templates from other merchants within the same audience group, selecting those that are most similar in product category and voucher type. When no suitable high-performing templates exist within the same merchant-audience group, RetrievalAgent searches cross-merchant examples to support rewriting. Each candidate message is represented by a dense embedding generated using a SentenceTransformer (Reimers and Gurevych, 2019) model all-MiniLM-L6-v2 Wang et al. (2020). Embeddings are computed based on structured metadata, including the product category hierarchy and voucher attributes. These vectors are indexed using Faiss Douze et al. (2024) for efficient similarity search. At inference time, the agent encodes the target message’s metadata into a dense vector and retrieves the top- most similar templates to serve as few-shot references for TemplateAgent. If both product and voucher metadata are missing, the system skips RetrievalAgent and falls back to rule-based rewriting.
3.3 TemplateAgent
@@ 4.1 Experimental Results
| Ori | Gen | (%) | Ori | Gen | (%) | Preferred | Rate (%) | |||
4.2 Ablation Studies
@@ 5 Conclusion
Limitations
While EvaluateAgent provides structured assessments of message quality, our evaluation is limited to offline settings. The generated templates have not yet been validated through online deployment or A/B testing, leaving real-world engagement impact unverified.

### v2

@@ Abstract.
1. Introduction
Customer Relationship Management (CRM) (Kumar and Reinartz, 2018) messages serve as a primary lever for user retention and revenue generation on modern e-commerce platforms. Merchants routinely push promotional notifications that announce new products, highlight price drops, or deliver loyalty rewards to segmented audiences through app and messaging channels. Crafting effective templates remains a bottleneck because most merchants lack dedicated marketing expertise, work under tight time constraints, and therefore rely on generic system-generated copy. Many CRM messages consequently underperform, leaving substantial engagement and conversion potential untapped.
Recent work has shown the effectiveness of LLM agents for structured decision-making in operational domains such as logistics (Ieva et al., 2025; Quan et al., 2025), supply chain management (Li et al., 2023; Quan and Liu, 2024b; Wang et al., 2025), and economic reasoning (Quan and Liu, 2024a; Guo and Yang, 2024; Liu and Quan, 2025a). To address this performance gap in CRM, we present CRMAgent (Figure 1), a multi-agent system built on large language models (LLMs) that automatically improves underperforming CRM message templates in e-commerce. Using historical engagement data, CRMAgent identifies low-performing messages, retrieves high-quality examples from similar contexts, and produces improved variants tailored to each merchant and audience segment group. The workflow is divided among four specialized agents: ContentAgent, RetrievalAgent, TemplateAgent, and EvaluateAgent, each responsible for a distinct subtask.
@@ 1. Introduction
2. Related Work
Recent multi-agent LLM systems adopt modular architectures to address structured generation tasks across diverse domains. For instance, Singh et al. (2024) design role-based agents for collaborative travel planning, while DocAgent (Yang et al., 2025) structures code documentation into sequential reader–writer–verifier stages. MAxPrototyper (Yuan et al., 2024) coordinates theme, text, and visual elements under a central controller for UI prototyping, and InvAgent (Quan and Liu, 2024b) demonstrates that agent specialization benefits inventory management in logistics. Similarly, CRMAgent applies a task-specialized multi-agent framework to CRM message rewriting, with agents dedicated to content analysis, retrieval, generation, and evaluation. Complementary to modularity, retrieval-augmented generation (RAG) improves LLM output by grounding in exemplar data or structured knowledge. Retrieve-and-Edit (Hashimoto et al., 2018), PEARL (Malik et al., 2024), and Chen et al. (2024) leverage retrieval to enhance preference modeling and factual accuracy. Recent work also explores temporal event modeling and retrieval (Liu and Quan, 2024, 2025b), combining LLMs with temporal point processes to unify event semantics and timing, while Liu et al. (2024) show that fine-tuning with retrieved exemplars improves personalization. Extending these directions, CRMAgent retrieves high-performing CRM templates based on audience, product, and voucher metadata, and generates tailored messages using marketing feedback.
3. Multi-Agent Collaboration for CRM Message Template Generation
@@ 3.1. ContentAgent
3.2. RetrievalAgent
RetrievalAgent retrieves high-quality CRM message templates from other merchants within the same audience group, selecting those that are most similar in product category and voucher type. When no suitable high-performing templates exist within the same merchant-audience group, RetrievalAgent searches cross-merchant examples to support rewriting. Each candidate message is represented by a dense embedding generated using a SentenceTransformer (Reimers and Gurevych, 2019) model all-MiniLM-L6-v2 (Wang et al., 2020). Embeddings are computed based on structured metadata, including the product category hierarchy and voucher attributes. These vectors are indexed using Faiss (Douze et al., 2024) for efficient similarity search. At inference time, the agent encodes the target message’s metadata into a dense vector and retrieves the top- most similar templates to serve as few-shot references for TemplateAgent. If both product and voucher metadata are missing, the system skips RetrievalAgent and falls back to rule-based rewriting.
3.3. TemplateAgent
@@ 4.1. Experimental Results
4.2. Ablation Studies